Use cylindrical coordinates to evaluate the integral. S SVO?-?? /o-+?=> p?dzaydx (a > 0) Enter the exact answer. S6 Soy Sa+=2Use cylindrical or spherical coordinates to evaluate the integral. 36—y2 2-x2y2 6* %* Son z? dz dx dy Enter the exact answer.
Answer #1

1) SSS. a 22 22 2² x²y2 x²d zdy dx +/2 a a²p2 o 0 0 p² cop?o. P d2 de doo s y=va²22 Z=0 O=0 त >X Projection onto xx plane P=0√72-x² 2 36-42 6. (3) ss 22 dz dady 0 x=86-12 22442 7/2 6 √72-p2 dz z? papdo 0=0 p=0 Z=p 6 312 I J L P (22-03) pol de = P. Z